OpenStudy (anonymous):

There are 2,100 bateria in a circular petri dish. The dish has radius of 40 milimeters. What is the approximate population density? A. 0.02 bacteria per square millimeter B. 0.42 bacteria per square millimeter C. 2.39 bacteria per square millimeter D. 52.5 bacteria per square millimeter

OpenStudy (hlares):

To find approximately how many bacteria there are per square millimeter, you first need to find the area of the circle. Since you are given the radius of 40 millimeters, then the area will be: A=(pi)*r^2=(pi)*40^2=5026.55 Then, since you are looking at bacteria per square millimeter, divide the number of bacteria by the area of the circle: 2,100/5026.55=0.417782... Thus, your answer will be B: 0.42 bacteria per square millimeter